# via Mixture-of-Curvature Experts

### v1

@@ 4.2 Hyperbolic Mixture-of-Curvature Module
Previous hyperbolic Transformer architectures fix each block to a single hyperbolic manifold, forcing the entire sequence to be embedded with a fixed curvature, restricting the flexibility of the hidden representations. Mixture-of-Experts (MoE) [26] used in Euclidean LLMs [10, 30, 13, 27, 9], where the model selectively activates only a subset of specialized "experts" for each token, enables LLMs to learn more diverse data patterns while remaining computationally efficient. However, the experts are still limited to one geometric space, restricting the collective granularity the experts can learn. Accommodating variable curvatures calls for a more flexible treatment: we propose the first Mixture-of-Curvature Experts (MiCE) module, where each expert operates on a distinct curvature space. Let $\mathbf{x}_{t}\in\mathbb{L}^{K_{1},n}$ be the $t$-th token input, then $\mathrm{MiCE}^{N_{s}}_{N_{r}}:\mathbf{x}_{t}\in\mathbb{L}^{K,n}\to\mathbf{x}_{t}\in\mathbb{L}^{K,n}$, where $N_{r}$ is the number of routed experts and $N_{s}$ is the number of shared experts. First, we pass $\mathbf{x}_{t}$ through a gating module to obtain the gating scores for each routed expert, denoted as $g_{t,i}$ for $1\leq N_{r}$, given as,
@@
The constants $\sqrt{K_{s,i}/K},\sqrt{K_{r,i}/K}$ project from the experts’ manifolds to the input manifold, ensuring that the output of each shared and routed expert lives on the same manifold. The outputs are then combined through the Lorentzian centroid [29], before performing a Lorentzian residual connection. Note that MiCE is indeed a valid hyperbolic module, which we expand on in Appendix B.
@@ 4.3 Efficiency via Hyperbolic Multi-Head Latent Attention
Previous hyperbolic Transformers for natural language applications focus mainly on hyperbolic self-attention, synonymous to naive dot-product attention mechanism in Euclidean LLMs. However, recent Euclidean LLMs have gradually adopted more efficient attention methods for enhanced scalability. The quadratic attention mechanism and the extant hyperbolic self-attention formulation suffer from heavy memory challenges for large-scale training. In this section, we propose hyperbolic Multi-Head Latent Attention (HMLA) to alleviate this bottleneck. Inspired by Euclidean MLA [11, 10], HMLA reduces the size of the KV cache during inference and the active memory consumption during training. We provide a high-level description of HMLA; detailed formulation can be found in Section B.2.
@@
Let $\mathbf{x}_{t}\in\mathbb{L}^{K,nh_{n}}$ be the $t$-token, where $n$ is the embedding dimension and $h_{n}$ is the number of heads. First, we project $\mathbf{x}_{t}$ via $\mathrm{HLT}$ to latent queries and key-value vectors, obtaining $\mathbf{c}_{t}^{Q},\mathbf{c}_{t}^{KV}$ of dimensions $n_{q},n_{kv}$ respectively such that $n_{q},n_{kv}\ll nh_{n}$. We then upward-project the latent vectors back to $h_{n}$ heads of dimension $n$ each via $\mathrm{HLT}$, obtaining $[\mathbf{k}_{t,i}^{C}]_{i\leq h_{n}},[\mathbf{v}_{t,i}^{C}]_{i\leq h_{n}}$ from $\mathbf{c}_{t}^{KV}$, and $[\mathbf{q}_{t,i}^{C}]_{i\leq h_{n}}$ from $\mathbf{c}_{t}^{Q}$. Then the projected keys are processed by positional encodings. Following previous works [11, 10], as RoPE is incompatible with MLA due to position coupling, we employ a decoupled HoPE scheme with HMLA. Through $\mathrm{HLT}$, we project $\mathbf{c}_{t}^{Q}$ to additional query vectors, $[\mathbf{q}_{t,i}^{R}]_{i\leq h_{n}}$, and $\mathbf{c}_{t}^{KV}$ to a shared key, $\mathbf{k}_{t}^{R}$, of dimension $n_{r}h_{n}$ and $n_{r}$ respectively ($n_{r}\ll nh_{n}$). The queries vectors, $\mathbf{q}_{t,i}^{C},\mathbf{q}_{t,i}^{R}$, and the key vectors, $\mathbf{k}^{C}_{t,i},\mathbf{k}^{R}_{t}$, are then concatenated together through Lorentzian concatenation [37, 48] for each $i$, where we obtain the final query and key vectors $\mathbf{q}_{t,i},\mathbf{k}_{t,i}$. The attention score and output are computed using Equation 3 with $[\mathbf{q}_{t,i}]_{t\leq h_{n}},[\mathbf{k}_{t,i}]_{t\leq h_{n}},[\mathbf{v}_{t,i}^{C}]_{t\leq h_{n}}$ as the queries, keys, and values, before concatenating all the heads together with Lorentzian concatenation. The concatenated result is passed through a final projection layer with $\mathrm{HLT}$.
@@
Memory complexity. During inference, HMLA only requires caching the latent key-value pairs. As a result, the memory footprint for HMLA is $O((n_{kv}+n_{r})L)$, where $L$ is the number of layers. In contrast, the hyperbolic self-attention used in previous hyperbolic Transformers (Equation 3) requires storing the full-sized keys and values, resulting in a memory complexity of $O(2nn_{h}L)$. By choosing $n_{kv},n_{r}\ll nn_{h}$, we have $(n_{kv}+n_{r})\ll 2nn_{h}$, resulting in a significantly smaller memory footprint while maintaining the same time complexity of $O((nn_{h})^{2})$. Additionally, the latent query projection also results in smaller active footprint during training. This collective mechanism enables far greater scalability.
@@ 4.4 Hyperbolic RMSNorm
Previous works have not implemented hyperbolic root mean square normalization (RMSNorm), which is widely used in popular Euclidean LLMs [18, 11, 10] due to its stability and robustness in both forward and backward passes. Here, we propose hyperbolic RMSNorm to fill the gap:
@@
Here, RMSNorm denotes the corresponding Euclidean operation. Equation 7 is equivalent to passing RMSNorm into the HRC functions from Hypformer [48]. This formulation is chosen over other methods of defining normalization layers through hyperbolic midpoint and tangent space operations [45, 3] for better numerical stability and computational efficiency.
@@ 4.5 Overall Architecture for Hyperbolic Large Language Models (HELM)
We introduce the framework for hyperbolic LLMs (HELM) based on the modules we introduced and developed in Section 4. In particular, we develop hyperbolic LLMs with Mixture-of-Curvature Experts (HELM-MiCE), a class of hyperbolic LLMs with MoE modules where each expert functions in a distinct curvature space. We also construct hyperbolic dense LLMs (HELM-D), which resembles classic decoder-only LLMs such as Llama [18].
The overall architecture is as follows (see Figure 3): tokenized text is first mapped to learned hyperbolic word embeddings. The resulting embeddings are passed through a series of hyperbolic decoder blocks, each consisting of two components: 1) the attention component, where the embeddings are normalized by a $\mathrm{RMSNorm}_{\mathcal{L}}$ layer, then processed by an attention block such as HMLA or self-attention, and finally added to the embeddings through Equation 2; and 2) the HFFN component, where the processed embeddings are again normalized by $\mathrm{RMSNorm}_{\mathcal{L}}$ before being passed through a HFFN block and residually added to the output of the attention block (Equation 2). For HELM-MiCE, the HFFN block can either be a dense block such as $\mathrm{HFFN}_{SG}$ or a MiCE block as defined in Section 4.2, where $\mathrm{HFNN}_{SG}$ is a hyperbolic SwiGLU FFN we built to be consistent with Euclidean LLMs (see Section B.3 for details). HELM-D contains only dense HFFN layers. The output of the final decoder block is then normalized once again before projected to logits for next-token prediction.
@@ 5 Experiments
We evaluate both HELM variants’ ability to answer MCQ questions in two popular benchmarks, MMLU [24] and ARC [8]. Additionally, we train an ablation HELM-MiCE with constant curvature across experts, comparing with HELM-MiCE models with varying curvature across experts.
@@ 5.1 Multichoice Benchmarking
We evaluate both HELM-MiCE and HELM-D at 100M-parameter scales, across a variety of benchmarks spanning STEM problem-solving, general knowledge, and commonsense reasoning. The dense models also serve as an ablation comparison with the MiCE models. We further scale the HELM-MiCE to 1B parameters as the smaller HELM-MiCE model outperformed HELM-D overall. Additional details regarding implementation and datasets can be found in Appendix B.
Table 1: Multichoice question-answering accuracy (%) of HELM models and their Euclidean counterparts. Bold denotes highest accuracy and underline denotes second-highest. HELM models consistently outperform their Euclidean counterparts, with HELM-MiCE achieving the highest accuracy overall.
@@
LLaMA 115M $\mathbf{21.1}$ $25.3$ $25.3$ $23.8$ $21.0$ $23.3$
HELM-D 115M $\underline{20.1}$ $25.9$ $27.0$ $\underline{25.8}$ $21.2$ $24.0$
[…]
HELM-MiCE 120M $19.3$ 26.026.0{26.0}26.0 $\underline{27.4}$ $24.7$ $\underline{23.5}$ $\underline{24.2}$
@@
HELM-MiCE 1B 19.819.8{19.8}19.8 $\mathbf{26.5}$ $\mathbf{28.4}$ 25.9 $\mathbf{23.7}$ $\mathbf{24.9}$
@@
Training Setup. We use the LLaMA3.1-8B tokenizer [18] for all models, with a vocabulary size of 128K. For HELM-D, we use hyperbolic self-attention and $\mathrm{HFFN}_{SG}$ for the decoder block. We use 6 heads and 6 layers for the 100M model. For HELM-MiCE, we use HMLA and a mixture of dense and MiCE layers, each with 2 active experts and one shared expert. We use 6 heads, 6 layers, and 4 experts per layer for the 100M model, and we use 14 heads, 16 layers, and 8 experts per layer for the 1B model. The experts have curvature initiated uniformly from $-0.1$ to $-2.0$. Additionally, we incorporate the auxiliary-loss-free load balancing scheme and complementary sequence-wise auxiliary loss from DeepSeekV3 [10] to encourage load balancing among the experts. Each model was trained on a cluster of 4 NVIDIA A6000 and 4 NVIDIA A800 GPUs with model and data parallelism, where at most 4 GPUs were used by each model.
@@
We use the English portion of the Wikipedia dataset [14] for training, comprising $\sim$6.4M rows of raw text, or roughly 5B tokens.
@@
Hyperbolic word embedding. Previous works [5, 39] directly map input tokens to trained hyperbolic embeddings. However, we experienced model instability when training the 1B models with this method. Therefore, we only train the space-like dimension of the Lorentz word embeddings.
@@
Baselines. We test against two popular Euclidean models: one dense model and one MoE model. For the dense model, we test HELM-D against LLaMA [18]. For the MoE model, we test HELM-MiCE against DeepSeekV3 [10]. We train both baselines from scratch at the same parameter scales as their HELM counterparts, with the same dataset, tokenizer, and training setup.
@@
Benchmarks. We evaluate on a variety of benchmarks, including STEM and general knowledge reasoning benchmarks such as MMLU [24], ARC-Challenging [8], and OpenbookQA [32], and commonsense reasoning benchmarks such as CommonsenseQA [41] ,such HellaSwag [50]. For MMLU and ARC, we use 5-shot predictions. For CommonsenseQA, OpenbookQA, and HellaSwag, we use 0-shot prediction.
@@
Results. The results are shown in Table 1. We report the accuracy of the models’ abilities to answer multiple choice questions from the benchmarks. We mainly focus on comparing models within the same architectural sub-family, i.e., dense models and MoE models are separately tested against each other. Both HELM variants consistently outperform their Euclidean counterparts. In particular, the smaller HELM-D model achieves higher accuracy than LLaMA on four out of the five benchmarks, whereas the smaller HELM-MiCE model outperforms the smaller DeepSeekV3 model on all five benchmarks. When comparing the $\sim$100M-scale HELM-D and HELM-MiCE models, the latter nearly always performs better and achieves higher accuracy overall. This reflects the effectiveness of using more flexible geometry, coupled with the overall fewer active parameters in HELM-MiCE. For the larger 1B-parameter models, HELM-MiCE consistently outperforms the 1B DeepSeekV3 model, achieving the highest accuracy overall.

### v2

@@ 4.2 Hyperbolic Mixture-of-Curvature Module
Previous hyperbolic Transformer architectures fix each block to a single hyperbolic manifold, forcing the entire sequence to be embedded with a fixed curvature, restricting the flexibility of the hidden representations. Mixture-of-Experts (MoE) [jacobs1991adaptive] used in Euclidean LLMs [deepseekai2024deepseekv3technicalreport, lepikhin2021gshard, fedus2022switch, jiang2024mixtral, dai2024deepseekmoe], where the model selectively activates only a subset of specialized "experts" for each token, enables LLMs to learn more diverse data patterns while remaining computationally efficient. However, the experts are still limited to one geometric space, restricting the collective granularity the experts can learn. Accommodating variable curvatures calls for a more flexible treatment: we propose the first Mixture-of-Curvature Experts (MiCE) module, where each expert operates on a distinct curvature space. Let $\mathbf{x}_{t}\in\mathbb{L}^{K_{1},n}$ be the $t$-th token input, then $\mathrm{MiCE}^{N_{s}}_{N_{r}}:\mathbf{x}_{t}\in\mathbb{L}^{K,n}\to\mathbf{x}_{t}\in\mathbb{L}^{K,n}$, where $N_{r}$ is the number of routed experts and $N_{s}$ is the number of shared experts. First, we pass $\mathbf{x}_{t}$ through a gating module to obtain the gating scores for each routed expert, denoted as $g_{t,i}$ for $1\leq N_{r}$, given as,
@@
The constants $\sqrt{K_{s,i}/K},\sqrt{K_{r,i}/K}$ project from the experts’ manifolds to the input manifold, ensuring that the output of each shared and routed expert lives on the same manifold. The outputs are then combined through the Lorentzian centroid [law2019lorentzian], before performing a Lorentzian residual connection. Note that MiCE is indeed a valid hyperbolic module, which we expand on in Appendix˜B.
@@ 4.3 Efficiency via Hyperbolic Multi-Head Latent Attention
Previous hyperbolic Transformers for natural language applications focus mainly on hyperbolic self-attention, synonymous to naive dot-product attention mechanism in Euclidean LLMs. However, recent Euclidean LLMs have gradually adopted more efficient attention methods for enhanced scalability. The quadratic attention mechanism and the extant hyperbolic self-attention formulation suffer from heavy memory challenges for large-scale training. In this section, we propose hyperbolic Multi-Head Latent Attention (HMLA) to alleviate this bottleneck. Inspired by Euclidean MLA [deepseekv2, deepseekai2024deepseekv3technicalreport], HMLA reduces the size of the KV cache during inference and the active memory consumption during training. We provide a high-level description of HMLA; detailed formulation can be found in Section˜B.2.
@@
Let $\mathbf{x}_{t}\in\mathbb{L}^{K,nh_{n}}$ be the $t$-token, where $n$ is the embedding dimension and $h_{n}$ is the number of heads. First, we project $\mathbf{x}_{t}$ via $\mathrm{HLT}$ to latent queries and key-value vectors, obtaining $\mathbf{c}_{t}^{Q},\mathbf{c}_{t}^{KV}$ of dimensions $n_{q},n_{kv}$ respectively such that $n_{q},n_{kv}\ll nh_{n}$. We then upward-project the latent vectors back to $h_{n}$ heads of dimension $n$ each via $\mathrm{HLT}$, obtaining $\mathbf{k}_{t,i}^{C}]_{i\leq h_{n}},[\mathbf{v}_{t,i}^{C}]_{i\leq h_{n}}$ from $\mathbf{c}_{t}^{KV}$, and $\mathbf{q}_{t,i}^{C}]_{i\leq h_{n}}$ from $\mathbf{c}_{t}^{Q}$. Then the projected keys are processed by positional encodings. Following previous works [deepseekv2, deepseekai2024deepseekv3technicalreport], as RoPE is incompatible with MLA due to position coupling, we employ a decoupled HoPE scheme with HMLA. Through $\mathrm{HLT}$, we project $\mathbf{c}_{t}^{Q}$ to additional query vectors, $\mathbf{q}_{t,i}^{R}]_{i\leq h_{n}}$, and $\mathbf{c}_{t}^{KV}$ to a shared key, $\mathbf{k}_{t}^{R}$, of dimension $n_{r}h_{n}$ and $n_{r}$ respectively ($n_{r}\ll nh_{n}$). The queries vectors, $\mathbf{q}_{t,i}^{C},\mathbf{q}_{t,i}^{R}$, and the key vectors, $\mathbf{k}^{C}_{t,i},\mathbf{k}^{R}_{t}$, are then concatenated together through Lorentzian concatenation [qu2022autoencoding, yang2024hypformer] for each $i$, where we obtain the final query and key vectors $\mathbf{q}_{t,i},\mathbf{k}_{t,i}$. The attention score and output are computed using Equation˜3 with $\mathbf{q}_{t,i}]_{t\leq h_{n}},[\mathbf{k}_{t,i}]_{t\leq h_{n}},[\mathbf{v}_{t,i}^{C}]_{t\leq h_{n}}$ as the queries, keys, and values, before concatenating all the heads together with Lorentzian concatenation. The concatenated result is passed through a final projection layer with $\mathrm{HLT}$.
@@
Memory complexity. During inference, HMLA only requires caching the latent key-value pairs. As a result, the memory footprint for HMLA is $O((n_{kv}+n_{r})L)$, where $L$ is the number of layers. In contrast, the hyperbolic self-attention used in previous hyperbolic Transformers (Equation˜3) requires storing the full-sized keys and values, resulting in a memory complexity of $O(2nn_{h}L)$. By choosing $n_{kv},n_{r}\ll nn_{h}$, we have $(n_{kv}+n_{r})\ll 2nn_{h}$, resulting in a significantly smaller memory footprint while maintaining the same time complexity of $O((nn_{h})^{2})$. Additionally, the latent query projection also results in smaller active footprint during training. This collective mechanism enables far greater scalability.
@@ 4.4 Hyperbolic RMSNorm
Previous works have not implemented hyperbolic root mean square normalization (RMSNorm), which is widely used in popular Euclidean LLMs [dubey2024llama3, deepseekv2, deepseekai2024deepseekv3technicalreport] due to its stability and robustness in both forward and backward passes. Here, we propose hyperbolic RMSNorm to fill the gap:
@@
Here, RMSNorm denotes the corresponding Euclidean operation. Equation˜7 is equivalent to passing RMSNorm into the HRC functions from Hypformer [yang2024hypformer]. This formulation is chosen over other methods of defining normalization layers through hyperbolic midpoint and tangent space operations [van2023poincar, Bdeir2024fully] for better numerical stability and computational efficiency.
@@ 4.5 Overall Architecture for Hyperbolic Large Language Models (HELM)
We introduce the framework for hyperbolic LLMs (HELM) based on the modules we introduced and developed in Section˜4. In particular, we develop hyperbolic LLMs with Mixture-of-Curvature Experts (HELM-MiCE), a class of hyperbolic LLMs with MoE modules where each expert functions in a distinct curvature space. We also construct hyperbolic dense LLMs (HELM-D), which resembles classic decoder-only LLMs such as Llama [dubey2024llama3].
The overall architecture is as follows (see Figure˜3): tokenized text is first mapped to learned hyperbolic word embeddings, which are then passed through a series of hyperbolic decoder blocks, each consisting of two components: 1) the attention component, where the embeddings are normalized by a $\mathrm{RMSNorm}_{\mathcal{L}}$ layer, then processed by an attention block such as HMLA or self-attention, and finally added to the embeddings through Equation˜2; and 2) the HFFN component, where the processed embeddings are again normalized by $\mathrm{RMSNorm}_{\mathcal{L}}$ before being passed through a HFFN block and residually added to the output of the attention block (Equation˜2). For HELM-MiCE, the HFFN block can either be a dense block such as $\mathrm{HFFN}_{SG}$ or a MiCE block as defined in Section˜4.2, where $\mathrm{HFNN}_{SG}$ is a hyperbolic SwiGLU FFN we built to be consistent with Euclidean LLMs (see Section˜B.3 for details). HELM-D contains only dense HFFN layers. The output of the final decoder block is then normalized once again before projected to logits for next-token prediction.
@@ 5 Experiments
We evaluate both HELM variants’ ability to answer MCQ questions in two popular benchmarks, MMLU [hendrycks2021measuring] and ARC [clark2018think]. Additionally, we train an ablation HELM-MiCE with constant curvature across experts, comparing with HELM-MiCE models with varying curvature across experts.
Runtime, Memory, and additional experiments. One common concern for hyperbolic models is that they could incur significant computational overhead when compared with their Euclidean counterparts. HELM models are within 1.55X in runtime and 1.11X in memory usage of the Euclidean counterparts for both the 100M and 1B models. Additional details are shown in Appendix˜D. We also show additional ablation and comparison with prior hyperbolic language models, for tasks such as machine translation, in Appendix˜D.
@@ 5.1 Multichoice Benchmarking
We evaluate both HELM-MiCE and HELM-D at 100M-parameter scales, across a variety of benchmarks spanning STEM problem-solving, general knowledge, and commonsense reasoning. The dense models also serve as an ablation comparison with the MiCE models. We further scale the HELM-MiCE to 1B parameters as the smaller HELM-MiCE model outperformed HELM-D overall. Additional details regarding implementation and datasets can be found in Appendix˜B.
Table 1: Multichoice question-answering accuracy (%) of HELM models and their Euclidean counterparts. Bold denotes highest accuracy and underline denotes second-highest. HELM models consistently outperform their Euclidean counterparts, with HELM-MiCE achieving the highest accuracy overall. The 100M model performances are average over 3 runs.
@@
LLaMA 115M $\mathbf{20.9}\pm 0.3$ $25.1\pm 0.3$ $25.4\pm 0.2$ $23.4\pm 0.5$ $21.0\pm 0.2$ $23.2\pm 0.2$
HELM-D 115M $\underline{20.3\pm 0.2}$ $25.9\pm 0.1$ $27.1\pm 0.4$ $\underline{25.6\pm 0.2}$ $\pm 0.3 24.1±124$$\pm 1$
DeepSeekV3 120M $19.3\pm 0.2$ $25.3\pm 0.1$ $24.0\pm 0.4$ $3.9\pm 0.3 22$$2\pm 0.3 22.$$2.2\pm 0.1$
HELM-MiCE 120M $19.7\pm 0.3$ $25.9\pm 0.2$} $\underline{27.7\pm 0.4}$ $24.4\pm 0.2$ $\underline{23.2\pm 0.5}$ $\underline{24.1\pm 0.1}$
@@
HELM-MiCE 1B 19.8{19.8} $\mathbf{26.5}$ $\mathbf{28.4}$ 25.9 $\mathbf{23.7}$ $\mathbf{24.9}$
@@
Training Setup. We use the LLaMA3.1-8B tokenizer [dubey2024llama3] for all models, with a vocabulary size of 128K. For HELM-D, we use hyperbolic self-attention and $\mathrm{HFFN}_{SG}$ for the decoder block. We use 6 heads and 6 layers for the 100M model. For HELM-MiCE, we use HMLA and a mixture of dense and MiCE layers, each with 2 active experts and one shared expert. We use 6 heads, 6 layers, and 4 experts per layer for the 100M model, and we use 14 heads, 16 layers, and 8 experts per layer for the 1B model. The experts have curvature initiated uniformly from 0.1-0.1 to 2.0-2.0. Additionally, we incorporate the auxiliary-loss-free load balancing scheme and complementary sequence-wise auxiliary loss from DeepSeekV3 [deepseekai2024deepseekv3technicalreport] to encourage load balancing among the experts. Each model was trained on a cluster of 4 NVIDIA A6000 and 4 NVIDIA A800 GPUs with model and data parallelism, where at most 4 GPUs were used by each model.
@@
We use the English portion of the Wikipedia dataset [wikidump] for training, comprising $\sim$6.4M rows of raw text, or roughly 5B tokens.
@@
Hyperbolic word embedding. Previous works [chen2021fully, HNN++] directly map input tokens to trained hyperbolic embeddings. However, we experienced model instability when training the 1B models with this method. Therefore, we only train the space-like dimension of the Lorentz word embeddings.
@@
Baselines. We test against two popular Euclidean models: one dense model and one MoE model. For the dense model, we test HELM-D against LLaMA [dubey2024llama3]. For the MoE model, we test HELM-MiCE against DeepSeekV3 [deepseekai2024deepseekv3technicalreport]. We train both baselines from scratch at the same parameter scales as their HELM counterparts, with the same dataset, tokenizer, and training setup.
@@
Benchmarks. We evaluate on a variety of benchmarks, including STEM and general knowledge reasoning benchmarks such as MMLU [hendrycks2021measuring], ARC-Challenging [clark2018think], and OpenbookQA [mihaylov2018openbookqa], and commonsense reasoning benchmarks such as CommonsenseQA [talmor2019commonsenseqa] ,such HellaSwag [zellers2019hellaswag]. For MMLU and ARC, we use 5-shot predictions. For CommonsenseQA, OpenbookQA, and HellaSwag, we use 0-shot prediction.
@@
Results. The results are shown in Table˜1. We report the accuracy of the models’ abilities to answer multiple choice questions from the benchmarks. We mainly focus on comparing models within the same architectural sub-family, i.e., dense models and MoE models are separately tested against each other. Both HELM variants consistently outperform their Euclidean counterparts. In particular, the smaller HELM-D model achieves higher accuracy than LLaMA on four out of the five benchmarks, whereas the smaller HELM-MiCE model outperforms the smaller DeepSeekV3 model on all five benchmarks. When comparing the $\sim$100M-scale HELM-D and HELM-MiCE models, the latter achieves comparable performance despite using overall significantly fewer active parameters. This reflects the effectiveness of using more flexible geometry. For the larger 1B-parameter models, HELM-MiCE consistently outperforms the 1B DeepSeekV3 model, achieving the highest accuracy overall. While we don’t provide standard deviation from multiple runs for the 1B models as is typical of models this size, we provide results of model performance across different stages of training in Appendix˜D and show that HELM-MiCE consistently outperform its Euclidean counterpart.